# Choose Your Diffusion: Efficient and flexible ways to accelerate the diffusion model in fast high energy physics simulation

### v1

@@ 1 Introduction
The Geant4 software toolkit [4] is applied for the detector simulation. The simulated primary particles would propagate through specific material and geometry, resulting vast amount of secondary particles produced in each propagated steps to achieve precision. As a consequence, the detector simulations for the modern high granularity detectors occupy the most computation resources. Reconstructing the physics objects like jets with complicated substructure for different physics process also plays a crucial role in providing the enough statistics for the data-driven analysis. In the upcoming High Luminosity LHC [5], the expected integrated luminosity will scale up to ten times the current level. It is nearly impossible to perform the full simulation for all the events with constrained computational budget, manifesting the necessity of fast simulation.
@@
One latest category of generative models, the Diffusion Model (DM) as well as its variations, has become increasingly dominant in various generative applications [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. There has been a few studies using DM in both fast detector and end-to-end event simulations.
@@
For end-to-end event simulation, there are a series of studies for generating the point cloud with DM [22, 23]. Notably, those studies also use architectures like transformer [22] or EPiC layer [23] that utilize the permutation invariance of jet constituents.
@@
We finally make our approaches portable by testing in different network and different datasets using the public fast and high-fidelity calorimeter dataset CaloChallenge [34] and the jet point cloud dataset JetNet [35]. The performance surpasses the previous model by several evaluation metrics.
@@ 2 Dynamics in the Diffusion Model
Figure 1: Upper: The schematic diagram for a noisy shower image (energy deposited in one layer of calorimeter) $x_{t}$ denoises along the ODE curvature with varying time steps to clean shower close to $x_{0}$. Lower: ODE and SDE ways to get the $x_{0}$, ODE flows along the certain direction toward $x_{0}$, where SDE will have extra noise injection in each time step before doing the ODE part. The Euler and Heun are the first and second order calculation for updating $x_{i}$.
@@
This determines the characteristics of the general ODE/SDE solver. ODE solvers often have smaller discretization errors than SDE solvers in smaller sampling steps. However, as the sampling steps become larger, the accumulated error for SDEs would be much smaller than that for ODEs, resulting in a smaller total error (better performance) in large sampling steps. The noise scheduler determines how fast those samplers should learn along different time steps. In the paper, we will cover up to seven mainstream samplers nowadays (four ODEs and three SDEs) over different schedulers.
@@
    • Variance Preserving schedulers used in [41], defined as a continous time step $t$ and $\sigma=\sqrt{(\exp{\frac{\beta_{\mathrm{max}}t^{2}}{2}+\beta_{\mathrm{min}}t}-1)}$.
    • The schedulers proposed by Karras [42]. $t_{i<N}=(\sigma_{\mathrm{max}}^{1/\rho}+\frac{i}{N-1}(\sigma_{\mathrm{min}}^{1/\rho}-\sigma_{\mathrm{max}}^{1/\rho}))^{\rho}$ and $\sigma_{N}=0$. It’s the monotonically decreasing function, so the noise added in a decreasing rate, $\rho$ determines the curvature. Larger $\rho$ means more steps near $\sigma_{\mathrm{min}}$ are shortened. Reasonable value for $\sigma_{\mathrm{min}},\sigma_{\mathrm{max}},\rho$ in our study should around $[0.01,0.001],[20,80],[5,10]$ respectively. This scheduler is helpful for generating good quality of data in low sampling step region.
    • Combine Karras schedulers proposed with Lu [43] by transforming the $\sigma$ to log scale. $t_{i<N}=(\log{(\sigma_{\mathrm{max}})}^{1/\rho}+\frac{i}{N-1}(\log{(\sigma_{\mathrm{min}})}^{1/\rho}-\log{(\sigma_{\mathrm{max}})}^{1/\rho}))^{\rho}$. The values for $\sigma_{\mathrm{min}},\sigma_{\mathrm{max}}$ remains the same region, $\rho=1$ to keep a healthy noise level range. This scheduler is even more decreasing than Karras schedulers, so the step size changes more rapidly in low sampling step while more slowly in last steps.
@@
    • The same DDPM samplers are used in [20, 39].
    • EDM Stochastic sampler with Heun’s method, as illustrated in Fig. 1, a noise injection step is performed to current time step $\hat{t_{i}}=t_{i}+\gamma t_{i}$ where $\gamma$ is determined by three parameters $\gamma=\min{(S_{churn}/N,\sqrt{2}-1)}$ when $S_{\mathrm{min}}\leq t_{i}\leq S_{\mathrm{max}}$. The $S_{\mathrm{max}},S_{\mathrm{min}}$ is the range for injection, $S_{churn}$ controls how fast it converges to zero noise. The optimal sets of this can make the sampling faster and better. Contrarily, a bad sets of those stochastic term will make possible degradation of the generated sample. In our study, such degradation only happens when extreme values are selected. The difference between this approach and Euler Maruyama is that the denoising steps in EDM samplers happens in some intermediate state $\hat{t_{i}}$, the difference become diminishing when N goes large. This SDE sampler, unlike other SDEs that only give satisfied result after a very large sampling steps, can converge quickly becuase of this.
@@
    • LMS sampler. For each iteration, the sampler will denoise the input data using the model and updates the $x_{i}$ using the linear multistep integration method with the calculated coefficients of predefined orders.
@@ 4.1 CaloChallenge
We choose to use the network from [20] which is composed of a conditional UNet with middle layer self-attention layer, cynlindircal convolution and Geometry Latent Mapping (GLaM). The latter two techniques are found to be beneficial for generating quantities in $r$ and $\alpha$ bins. In order to demostrate the portability of our studies, the neural network architect remains unchanged from the version in the literature [20].
@@ 4.2 JetNet
For JetNet dataset, we follow the preprocessing procedure from [22, 23]. The Equivariant Point Cloud (EPiC) layer based on the deep sets framework is chosen to be used as it shows great quality of generated classes while having lighter architecture than transformer-based models [23].
@@
We follow the way from [42] to rewrite the score function to a denoiser function which minimizing the $L2$ denoising error at each noise level:
$\nabla_{x}\log{p_{t}(x)}=(D(x)-x)/t^{2}$ (8)
A set of preconditioning parameters used to predict noise $\epsilon$ with this denoiser function:
@@
where $c_{in},c_{out}$ maps the input and output data magnitude, and $c_{\mathrm{skip}}$ controls the skip connection of the network. $c_{in}=1/\sqrt{t^{2}+\sigma_{c}^{2}}$ and $c_{out}=(t*\sigma_{c})/\sqrt{t^{2}+\sigma_{c}^{2}}$, $\sigma_{c}$ is the constant value, a reasonable value is around $[0.5,1.5]$ to avoid large variation for gradient scale. The $c_{\mathrm{skip}}=\sigma_{c}^{2}/(t^{2}+\sigma_{c}^{2})$ is important quantity to partially mask the skip connection when the noise level goes large. This can prevent the network from inaccurate prediction. The setting is important in our case given noise varies widely in most of samplers we show later. When the $t$ goes higher, the $c_{\mathrm{skip}}$ becomes smaller. In a way, this compels the skip connection to be deactivated, encouraging the network to prioritize learning the signal for overriding the input rather than predicting noise to correct it. We choose the continuous timestep training where $\log{t}=\mathcal{N}(P_{\mathrm{mean}},P^{2}_{\mathrm{std}})$ where $P_{\mathrm{mean}},P_{\mathrm{std}}$ is -1.2 and 1.2 correspondingly.
From our study, the choice of the weighting function is very important as it not only determines the speed of loss convergence but also the quality of the outputs and how the specific energy distribution tends to evolve across different time steps. To mitigate the instability of the optimization over the training, and balance the gradient magnitude over diverse noise levels, one loss weight category is often mentioned $-$ Signal to noise ratio (SNR) weighting [29, 48, 49, 50]. We choose the min-SNR weighting as it is found to be beneficial for training high-resolutions dataset. The original min-SNR weighting proposed from [51] defined to be min{SNR(t),$\gamma$}, where $\gamma$ is the upper limit for weighting value, usually set to be 2-5. This could avoid the model putting too much weight on the low noise level.
@@
$w(t)=(t*\sigma_{c})^{2}/(t^{2}+\sigma_{c}^{k})^{2}$ (10)
that $t$ and $\sigma_{c}$ is the same as the preconditioning parameters, the reasonable value of k in the range of $[2,4]$. This function is even with one minimum value of 0 around zero $t$ and two local maxima on positive and negative t-axes. The constant k determines the local maximum weight value, and the SNR value where the maxima start to smoothly fall down to 0. The advantages of this weighting scheme will be shown later in Sec 5 with comparison of default constant (unit) weighting strategy.
Figure 2: The distribution for total shower energy divided by the incident particle energy, tale dotted line is the GEANT4 reference, red solid line is the generated distribution by 79 steps EDM samplers. Upper: 30, 70, 100, 200 epochs default denoiser training with unit weight. Lower: 30, 70, 100, 200 epochs training with our proposed min-SNR weighting. The default training still struggles to learn this variable after 200 epochs whereas the training with proposed weighting strategy already gives good shape alignment after 100 epochs.
@@ 5 Performance
We have previously discussed various schemes for the backward diffusion process. In order to comprehensively compare their performance and enhance understanding of how to choose the appropriate samplers and schedulers for a specific purpose, we offer several methods for evaluation.
@@ 5.1 CaloChallenge
There are multiple important distributions related to the layer and voxel energies for CaloChallenge dataset. We compare the generated samples with GEANT4 as reference. The voxel energy distribution defined as the distribution of raw generated low level voxel energies. The total energy distribution defined as the distribution of total energy of the generated shower. The center energy of the shower defined as $\bar{x}=\frac{\langle x_{i}E_{i}\rangle}{\sum E_{i}}$ for cell location $x_{i}$ and energy $E_{i}$, and shower width defined as $\sqrt{\bar{x^{2}}-\bar{x}^{2}}$, we will show these two quantities at the polar coordinate $\alpha$ and $r$. The mean layer energy distribution defined as the mean of generated layer energies. The shower response $E_{\mathrm{ratio}}$ defined as the total shower energy divided by the incident particle energies, this quantity is extremely important as a correction factor for energy calibration of many particles at LHC. We will also evaluate the separation power between generated $E_{\mathrm{ratio}}$ and reference one.
@@
In addition, a DNN classifier with 2 layers, each comprising 2048 nodes and a 0.2 dropout rate, is employed to assess the area under the receiver operating characteristic curve (AUC) for all high-level features, including center energies, shower width, and layer energies, of both generated samples and references. An AUC value that close to 0.5 suggests that the generated and target classes are nearly indistinguishable. Conversely, a higher AUC value indicates a more evident distinction between the quantities of the two classes.
@@
Figure 3: Comparison plots for center energy of 5000 showers in radial bins. left: EDM with 30 steps, middle: EDM with 79 steps, right: Restart with 30 steps.
First, a comparison plots for $E_{\mathrm{ratio}}$ with default EDM training loss weight and our proposed training weight shown in Fig. 2. Several observations can be made from the plots. The choice of the weighting function influences how the kinematics distribution evolves over time. After 200 training epochs, the model trained with the default weighting still struggles to learn the $E_{\mathrm{ratio}}$ distribution. In contrast, employing proposed min-SNR weighting results in the generated $E_{\mathrm{ratio}}$ aligning with the reference shape after only 100 epochs. This observation affirms that min-SNR weighting expedites the convergence of high-level features to achieve better performance compared to constant weighting.  [^2] A 1D layer diffusion can generate better $E_{\mathrm{ratio}}$ distribution. But it might be over simplified as the model only considers the layer deposited energy. The paper only considers the low level voxel diffusion as a more realistic and challenging case for fast calorimeter simulation
Few plots for center energy in $r$ direction are selected in Fig. 3. Notably, the EDM samplers and Restart samplers with Karras schedulers can learn this quantities really well. The ratio difference is almost indistinguishable with naked eyes. We can tell from the small ($\sim$30) steps EDM, the differences in last three radial bins are larger meaning it underestimates the lower energy region. When the steps become larger, we see the EDM predicts well in all radial bins. The Restart sampler can achieve this with only 30 steps. Because the Restart samplers works with a predefined parameters for configuring the EDM ODE. The generation time is directly proportional with the restart iterations in the configuration, usually longer than default EDM samplers. A lack of principled way for these hyperparameters is the inherent limitation for this samplers. Later we will see the best performance is achieved around 30-70 steps as we only uses specific configurations in this range.
Then the mean layer energies, total shower energies and voxel energies for the same three cases are shown in Fig. 456 (right).
@@
Figure 5: Comparison histograms for 5000 total shower energies. left: EDM with 30 steps, middle: EDM with 79 steps, right: Restart with 30 steps.
@@
Figure 6: Comparison histograms for raw voxel energies of 5000 showers: Left: 30 steps EDM, 79 steps EDM, 30 steps Restart samplers; Right: 400 steps DDPM, 36 steps LMS, and 15 steps DPM-Solver $++$ samplers.
@@
Most of the samplers need a relatively long sampling time to learn the mean energy of the last few layer correctly since they have voxels with lowest energy thresholds. The low steps EDM sampler has good performance on low energy voxel while mediocre performance on high energy tails of voxel distributions and low energy shower. DPM-solver samplers with Karras scheduler does very well on predicting the voxel energy around 15 steps, achieving an exceptional agreement on both low and high energy tails shown in the left plot of Fig. 6. All proposed samplers have comparable and better performance with smaller steps. Indeed some of them are struggling to match the low voxel energy, the presence of the tail is probably a consequence of model itself and too low energy threshold. It is expected this deficiency should not have big impact on the downstream reconstruction of shower.
However, DPM-solver sampler seems to perform poorly on predicting the low and high-level features simultaneously. The Uni-PC sampler has very similar performance with DPM-solver sampler and only gives slightly better performance over very low sampling (5-10) steps. The lowest AUC for high level feature classifiers is around 60 at 20th step. We showed the AUC plots for evaluating different samplers 20 times in Fig. 7. The EDM and Restart samplers surpasses the previous 400 steps DDPM samplers at 50, 25 steps correspondingly, combined with other plots we showed, achieving the same and better performance with a 3-8 times speedup. The LMS ODE has a better performance than EDM below 45 steps, but reaches a worse convergence than other samplers. The best AUC value for Restart samplers achieves at 30-36 and 79 steps, this might because we use the predefined configuration specifically for those steps. Most of the samplers we use actually have more parameters to tune than the baseline DDPM. For the common ODE samplers, a closer matching noise level during sampling would often yield better results., LMS sampler involves an additional parameter "order" of the coefficient which makes the generation time longer as it increases. A possible future development is to fine-tune the configuration in higher steps.
Figure 7: AUC values for different samplers as a function of sampling steps, the shaded area is possible AUC value range for 20 evaluations through independent classifier training.

### v2

@@ 1 Introduction
The Geant4 software toolkit [4] is applied for the detector simulation. The simulated primary particles would propagate through specific material and geometry, resulting vast amount of secondary particles produced in each propagated steps to achieve precision. As a consequence, the detector simulations for the modern high granularity detectors occupy the most computation resources[5]. Reconstructing the physics objects like jets with complicated substructure for different physics process also plays a crucial role in providing the enough statistics for the data-driven analysis. In the upcoming High Luminosity LHC [6], the expected integrated luminosity will scale up to ten times the current level. It is nearly impossible to perform the full simulation for all the events with constrained computational budget, manifesting the necessity of fast simulation.
@@
One latest category of generative models, the Diffusion Model (DM) as well as its variations, has become increasingly dominant in various generative applications [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. There has been a few studies using DM in both fast detector and end-to-end event simulations.
@@
For end-to-end event simulation[20, 21], there are a series of studies for generating the point cloud with DM [25, 26]. Notably, those studies also use architectures like transformer [25] or EPiC layer [26] that utilize the permutation invariance of jet constituents.
@@
We finally make our approaches portable by testing in different network and different datasets using the public fast and high-fidelity calorimeter dataset CaloChallenge [37, 38] and the jet point cloud dataset JetNet [39]. The performance surpasses the previous model by several evaluation metrics[37, 8, 25, 38, 40].
@@ 2 Dynamics in the Diffusion Model
Figure 1: Upper: The schematic diagram for a noisy shower image (energy deposited in one layer of calorimeter) $x_{t}$ denoises along the ODE curvature with varying time steps to clean shower close to $x_{0}$. Lower: Arrows indicate a schematic sketch how ODE and SDE samplers reach the $x_{0}$ along timesteps, ODE flows along the certain direction toward $x_{0}$. $t_{i}$ is the $i^{th}$ timestep, SDE will have extra noise injection $\gamma t_{i}$ in each time step before moving to next step. The Euler and Heun are the first and second order calculation for updating $x_{i}$.
@@
This determines the characteristics of the general ODE/SDE solver. ODE solvers often have smaller discretization errors than SDE solvers in smaller sampling steps[48, 49]. However, as the sampling steps become larger, the accumulated error for SDEs would be much smaller than that for ODEs, resulting in a smaller total error (better performance) in large sampling steps. The noise scheduler determines how fast those samplers should learn along different time steps. In the paper, we will cover up to seven mainstream samplers nowadays (four ODEs and three SDEs) over different schedulers.
@@
    • Variance Preserving schedulers as a more generalized version of cosine beta schedulers used in both score matching [46] and DDPM, defined as a continous time step $t$ and $\sigma=\sqrt{(\exp{\frac{\beta_{\mathrm{max}}t^{2}}{2}+\beta_{\mathrm{min}}t}-1)}$ where $\beta_{\mathrm{min}}$ and $\beta_{\mathrm{max}}$ are constant to determine the lower and upper bound of noise level.
    • The schedulers proposed by Karras [47]. $\sigma_{i<N}=(\sigma_{\mathrm{max}}^{1/\rho}+\frac{i}{N-1}(\sigma_{\mathrm{min}}^{1/\rho}-\sigma_{\mathrm{max}}^{1/\rho}))^{\rho}$ and $\sigma_{N}=0$. It’s the monotonically decreasing function, so the noise added in a decreasing rate, $\rho$ determines the curvature. Larger $\rho$ means more steps near $\sigma_{\mathrm{min}}$ are shortened. Reasonable value for $\sigma_{\mathrm{min}},\sigma_{\mathrm{max}},\rho$ in our study should around $[0.01,0.001],[20,80],[5,10]$ respectively. This scheduler is helpful for generating good quality of data in low sampling step region.
    • Combine Karras schedulers proposed with Lu [50] by transforming the $\sigma$ to log scale. $\sigma_{i<N}=(\log{(\sigma_{\mathrm{max}})}^{1/\rho}+\frac{i}{N-1}(\log{(\sigma_{\mathrm{min}})}^{1/\rho}-\log{(\sigma_{\mathrm{max}})}^{1/\rho}))^{\rho}$. The values for $\sigma_{\mathrm{min}},\sigma_{\mathrm{max}}$ remains the same region, $\rho=1$ to keep a healthy noise level range. This scheduler is even more decreasing than Karras schedulers, so the step size changes more rapidly in low sampling step while more slowly in last steps.
@@
    • The same DDPM samplers are used in [23, 44]. The Euler-Maruyama method approximates the reverse diffusion process by iteratively denoising the sample while adding stochastic noise at each discrete timestep.
    • EDM Stochastic sampler with Heun’s method [47], as illustrated in Fig. 1, a noise injection step is performed to current time step $\hat{t_{i}}=t_{i}+\gamma t_{i}$ where $\gamma$ is determined by three parameters $\gamma=\min{(S_{churn}/N,\sqrt{2}-1)}$ when $S_{\mathrm{min}}\leq t_{i}\leq S_{\mathrm{max}}$. The $S_{\mathrm{max}},S_{\mathrm{min}}$ is the range for injection, $S_{churn}$ is the stochasticity strength which give the lower bound to start adding noise. It controls how fast the sampler converges to zero noise. The optimal sets of parameters can make the sampling faster and better. Contrarily, a bad sets of those stochastic term will make possible degradation of the generated sample. In our study, such degradation only happens when extreme values (e.g. large $\gamma$, or $S_{churn}$ close to initial timestep) are selected. The difference between this approach and Euler Maruyama is that the denoising steps in EDM samplers happens in some intermediate state $\hat{t_{i}}$ instead of actual timestep state $t$, the difference between two states become diminishing when N goes large. This SDE sampler, unlike other SDEs that only give satisfied result after a very large sampling steps, can converge quickly if one can manage to tune those stochastic parameters well.
@@
    • LMS sampler[51]. For each iteration, the sampler will denoise the input data using the model and updates the $x_{i}$ using the linear multistep integration method with the calculated coefficients of predefined orders.
@@ 4.1 CaloChallenge
We choose to use the network from [23] which is composed of a conditional UNet with middle layer self-attention layer, cynlindircal convolution and Geometry Latent Mapping (GLaM). The latter two techniques are found to be beneficial for generating quantities in $r$ and $\alpha$ bins. In order to demostrate the portability of our studies, the neural network architect remains unchanged from the version in the literature [23]. The original training is tailored for discrete noise level samplers like DDPM [44].We have integrated our training setups, including different loss function designs and data preprocessing, into the new training process. We use publicly available pre-trained models with 400 DDPM steps as the baseline to compare the performance.
@@ 4.2 JetNet
For JetNet dataset, we follow the preprocessing procedure from [25, 26]. The Equivariant Point Cloud (EPiC) layer based on the deep sets framework is chosen to be used as it shows great quality of generated classes while having lighter architecture than transformer-based models [26]. With the same training procedure changes, we further modify the architectures with similar sinusoidal and cosine positional embedding by leveraging sine and cosine functions of the noise levels.
@@
4.3 Training Details
We follow the way from [47] to rewrite the score function to a denoiser function which minimizing the $L^{2}$ denoising error between $x$ and model output $D(x)$ at each noise level:
$\nabla_{x}\log{p_{\sigma}(x)}=(D(x)-x)/\sigma^{2}$ (8)
If the network is preconditioned with $\sigma$ independent skip connections, $D$ can be written as $D_{\theta}(x,\sigma)=c_{\mathrm{skip}}(\sigma)x+c_{out}(\sigma)F_{\theta}(c_{in}(\sigma)x,\sigma,z)$. A set of preconditioning parameters used to predict noise $\epsilon$ with this denoiser function:
@@ 4.3 Training Details
$=\mathbb{E}_{\sigma,\epsilon}[w(\sigma)||F_{\theta}(c_{in}*x_{\sigma},\sigma,z)-\frac{1}{c_{out}}(x_{0}-c_{\mathrm{skip}}(\sigma)*(x_{\sigma}))||^{2}_{2}]$
where $F_{\theta}$ is the preconditioned network, $c_{in},c_{out}$ maps the input and output data magnitude, and $c_{\mathrm{skip}}$ controls the skip connection of the network. The parameters $c_{in}=1/\sqrt{\sigma^{2}+\sigma_{c}^{2}}$ and $c_{out}=(\sigma*\sigma_{c})/\sqrt{t^{2}+\sigma_{c}^{2}}$, $\sigma$ is the current noise levels, $\sigma_{c}$ is the constant value, a reasonable value is around $[0.5,1.5]$ to avoid large variation for gradient scale. $c_{\mathrm{skip}}=\sigma_{c}^{2}/(\sigma^{2}+\sigma_{c}^{2})$ is important quantity to partially mask the skip connection when the noise level goes large. This can prevent the network from inaccurate prediction. The setting is important in our case given noise varies widely in most of samplers we show later. When the $\sigma$ goes higher, the $c_{\mathrm{skip}}$ becomes smaller. In a way, this compels the skip connection to be deactivated, encouraging the network to prioritize learning the signal for overriding the input rather than predicting noise to correct it. We choose the continuous timestep training where $\log{t}=\mathcal{N}(P_{\mathrm{mean}},P^{2}_{\mathrm{std}})$ where $P_{\mathrm{mean}},P_{\mathrm{std}}$ is -1.2 and 1.2 correspondingly.
From our study, the choice of the weighting function is very important as it not only determines the speed of loss convergence but also the quality of the outputs and how the specific energy distribution tends to evolve across different time steps. To mitigate the instability of the optimization over the training, and balance the gradient magnitude over diverse noise levels, one loss weight category is often mentioned $-$ Signal to noise ratio (SNR) weighting [32, 54, 55, 56]. We choose the min-SNR weighting as it is found to be beneficial for training high-resolutions dataset. The original min-SNR weighting proposed from [57] defined to be min{SNR(t),$\gamma$}, where SNR(t) is defined as $\frac{1-\sigma^{2}}{\sigma^{2}}$, $\gamma$ is the upper limit for weighting value, usually set to be 2-5. This could avoid the model putting too much weight on the low noise level.
@@
$w(t)=(\sigma*\sigma_{c})^{2}/(\sigma^{2}+\sigma_{c}^{k})^{2}$ (10)
that $\sigma$ and $\sigma_{c}$ are the same as the preconditioning parameters, the reasonable value of k in the range of $[2,4]$. This function has one minimum value of 0 around zero $t$ and two local maxima on positive and negative t-axes. The constant k determines the local maximum weight value, and the SNR value where the maxima start to smoothly fall down to 0. The advantages of this weighting scheme will be shown later in Sec 5 with comparison of default constant (unit) weighting strategy.
Figure 2: The distribution for total shower energy divided by the incident particle energy, tale shaded region is the GEANT4 reference, the solid line is the generated distribution by 79 steps EDM samplers with different weight at 30, 50, 70, 100, 200 epochs. Purple: default denoiser training with unit weight. Pink: proposed min-SNR weighting. The default training still struggles to learn this variable after 200 epochs whereas the training with proposed weighting strategy already gives good shape alignment after 100 epochs.
@@ 5 Performance
Various schemes for the backward diffusion process have been discussed in previous section. To comprehensively compare their performance and enhance understanding of choosing the appropriate samplers and schedulers for a specific purpose, we offer several methods for evaluation.
@@ 5.1 CaloChallenge
There are multiple important distributions related to the layer and voxel energies for CaloChallenge dataset. We compare the generated samples with GEANT4 as reference. The voxel energy distribution is defined as the distribution of raw generated low level voxel energies. The total energy distribution is defined as the distribution of total energy of the generated shower. The center energy of the shower is defined as $\bar{x}=\frac{\langle x_{i}E_{i}\rangle}{\sum E_{i}}$ for cell location $x_{i}$ and energy $E_{i}$, and shower width is defined as $\sqrt{\bar{x^{2}}-\bar{x}^{2}}$, we will show these two quantities at the polar coordinate $\alpha$ and $r$. The mean layer energy distribution defined as the mean of generated layer energies. The shower response $E_{\mathrm{ratio}}$ is the total shower energy divided by the incident particle energies, this quantity is extremely important for energy calibration of many particles at LHC. As it will be used to correct the deposited shower energy inside calorimeter into the real particle energy, the small difference between generated sample and GEANT4 means more consistency for particle reconstruction procedure of both fast and full simulation. We will also evaluate the separation power between generated $E_{\mathrm{ratio}}$ and reference one.
@@
In addition, a public evaluation network [37] DNN classifier with 2 layers, each comprising 2048 nodes and a 0.2 dropout rate, is employed to assess the area under the receiver operating characteristic curve (AUC) for all high-level features, including center energies, shower width, and layer energies, of both generated samples and references. An AUC value that close to 0.5 suggests that the generated and target classes are nearly indistinguishable. Conversely, a higher AUC value indicates a more evident distinction between the quantities of the two classes.
@@
Figure 3: Comparison plots for average center energy of 5000 showers in radial bins. left: EDM with 30 steps, middle: EDM with 79 steps, right: Restart with 30 steps. Those samplers are chosen because of their good sample quality and inference speed trade-off, while the difference almost indistinguishable.
@@
First, a comparison plot for $E_{\mathrm{ratio}}$ with default EDM training loss weight and our proposed training weight is shown in Fig. 2. The choice of the weighting function influences how the kinematics distribution evolves over time. After 200 training epochs, the model trained with the default weighting still struggles to learn the $E_{\mathrm{ratio}}$ distribution. In contrast, employing proposed min-SNR weighting results in a better alignments with reference histogram, the generated $E_{\mathrm{ratio}}$ aligning with GEANT4 sample after only 100 epochs. This observation affirms that min-SNR weighting expedites the convergence of high-level features to achieve better performance compared to constant weighting.  [^2] A 1D layer diffusion can generate better $E_{\mathrm{ratio}}$ distribution. But it might be over simplified as the model only considers the layer deposited energy. The paper only considers the low level voxel diffusion as a more realistic and challenging case for fast calorimeter simulation
Few plots for center energy in $r$ direction are selected in Fig. 3. We choose the EDM Heun SDE and Restart samplers that have good sample quality and inference speed trade-off. Notably, the EDM samplers and Restart samplers with Karras schedulers can learn this quantities really well. The ratio difference is almost indistinguishable with naked eyes. The default scheduler for the EDM sampler is Karras. To better illustrate the impact of different noise schedulers on the same sampler, a broader exploration of scheduler choices—such as Karras, DPM++ , and Lu schedulers with varying decay rates for $\sigma$ is conducted. For small ($\sim$30) steps EDM, the differences in last three radial bins are larger meaning it underestimates the lower energy region. When the steps become larger, we see the EDM predicts well in all radial bins. The Restart sampler can achieve this with only 30 steps. Given the Restart samplers works with a predefined parameters for configuring the EDM ODE, the generation time is directly proportional with the restart iterations in the configuration, usually longer than default EDM samplers. A lack of principled way for these hyperparameters is the inherent limitation for this sampler. Later we will see the best performance is achieved around 30-70 steps as we only use specific configurations in this range.
Then the mean layer energies, total shower energies, shower response, and voxel energies for the same three cases are shown in Fig. 456.
@@
Figure 5: Comparison histograms for 100k shower energies. left: shower response, right: total shower energy from 400 steps DDPM (CaloDiffusion), 79 steps EDM, 30 steps Restart 36 steps LMS, 15 steps DPM-Solver ++, and 15 steps UniPC samplers
Figure 6: Comparison histograms for raw voxel energies of 100k showers: Left: 30 steps EDM, 79 steps EDM, 30 steps Restart samplers; Right: 400 steps DDPM (CaloDiffusion), 36 steps LMS, and 15 steps DPM-Solver $++$ samplers. Bottom is the percentage of difference between Geant4 and generated samples. The upward arrows indicate that the differences fall beyond the window at certain voxel energy levels.
@@
The low steps EDM sampler has good performance on low energy voxel while mediocre performance on high energy tails of voxel distributions. In Fig. 5, we have seen the shape of shower response between generated and reference sample for Restart and EDM samplers are more aligned than other samplers, especially around the peak of histogram ($\sim$0.8). DPM-solver sampler with Karras scheduler does very well on predicting the voxel energy around 15 steps, achieving an exceptional agreement on both low and high energy tails shown in the left plot of Fig. 6. The energy thresholds also affect low voxel energies generated by different samplers. For the baseline comparison, we used the pre-trained CaloDiffusion [23] model for dataset 2 with default 400 steps DDPM has 4-10% difference in the intermediate energy level (5-1000 MeV), and less than 30% difference with GEANT4 until 0.7 MeV. A clear difference has been seen below 0.7 MeV. The same scenario happens to EDM, Restart and LMS samplers, but all three samplers show more robust performance (i.e. a smaller excess in low energy tail). Generated showers produced by EDM and Restart samplers both have much less difference with GEANT4 around the intermediate energy level. Two new ODE samplers, DPM-Solver and UniPC, hold only less than 10% difference in low energy region, nevertheless perform worse on other regions. Indeed, some of them are struggling to match the low voxel energy, the presence of the tail is probably a consequence of model itself and too low energy threshold. It is expected this deficiency should not have big impact on the downstream reconstruction of shower.
Figure 7: AUC values for different samplers as a function of evaluation steps, the shaded area is possible AUC value range for 20 evaluations through independent classifier training. The effective ratio is defined as the ratio between actual AUC from generated sample with different solvers minus the best possible AUC value 0.5, divided by the AUC value from generated sample with DDPM minus 0.5 $\frac{\text{AUC}_{gen}-0.5}{\text{AUC}_{DDPM}-0.5}$ Lower is better, value below 1 means better performance than DDPM at certain evaluation steps.